# The Endoscapes Dataset for Surgical Scene Segmentation, Object Detection, and Critical View of Safety Assessment: Official Splits and Benchmark

### v1

@@ 1 Introduction
Automatic assessment of the critical view of safety (CVS) is an important problem in surgical data science that has been gaining interest in recent years Strasberg et al (1995); Mascagni et al (2020, 2021b); Murali et al (2023b); Ban et al (2023); Ríos et al (2023); Murali et al (2023a). Unlike many surgical data science tasks that have been studied thus far (e.g. phase recognition, tool detection/segmentation), CVS assessment is particularly challenging because it relies on accurate and precise identification of fine-grained anatomical structures and concepts. To enable research into this task, we present Endoscapes, a dataset of laparoscopic cholecystectomy videos with annotations specifically targeted at tackling automated CVS assessment: namely, frame-level CVS annotations as well as segmentation/bounding boxes of critical anatomical structures/regions relevant for CVS assessment (such as the hepatocystic triangle, cystic artery/duct, and cystic plate).
In this report, we begin by describing the development of this dataset, which represents the culmination of several works Mascagni et al (2021b); Alapatt et al (2021); Murali et al (2023b). Then, we provide a detailed enumeration of the contents of Endoscapes2023, released alongside this report, which comprises three sub-datasets categorized by annotation type and number of videos: Endoscapes-CVS201, Endoscapes-BBox201, and Endoscapes-Seg50. Finally, we present benchmarks for three different tasks (object detection, instance segmentation, and CVS prediction).
@@ 2 Context
Automated CVS assessment was first tackled by our prior work DeepCVS Mascagni et al (2021b), where we present a method for joint anatomical segmentation and CVS prediction. For this initial work, we hand-picked 2854 images where we deemed CVS assessable (in the anterior view) from 201 laparoscopic cholecystectomy (LC) videos, then annotated these images with each of the three CVS criteria (three binary image-level annotations), based on the protocol established in Mascagni et al (2020). We also selected a subset of 402 images ($\sim$20%), balancing optimal CVS achievement (3/3 criteria achieved) and suboptimal CVS achievement (0-2/3 criteria achieved) and further annotated them with segmentation masks. This initial dataset, while instrumental in enabling a proof-of-concept study, was characterized by two critical limitations: (1) the frame selection process introduces bias during both training and evaluation, and (2) the dataset is very small, with a test set of only 571 images for CVS assessment and 80 images for segmentation.
Recognizing these issues, in a follow-up work Alapatt et al (2021), we collected a vastly expanded dataset from the same 201 videos. Here, rather than hand-selecting frames from the videos, we simply extract frames at 1fps from the region of video where CVS is adjudged to be evaluable [^4] This region, which is a subset of the dissection phase, begins when “any of the 3 CVS criteria is considered evaluable” and ends “when the first clip is applied on the cystic duct or the cystic artery”, as defined in Mascagni et al (2021a)., thereby eliminating selection bias. For these frames, we then annotate segmentation (5 anatomical structures and 1 tool class) and CVS at regular intervals: 1 frame every 30 seconds for segmentation Alapatt et al (2021) and 1 frame every 5 seconds for CVS Murali et al (2023b). Removing the manual frame selection process introduced numerous issues during annotation, to which end we developed a more robust annotation protocol for both segmentation and CVS assessment Mascagni et al (2021a). Finally, following this updated protocol and incorporating several rounds of review to ensure optimal consistency, we annotated 1933 total images with segmentation masks and 11090 images with CVS annotations from the 201 LC videos. The CVS annotations in particular were done independently by three expert clinicians to reduce bias; for evaluation, we define the ground-truth as the consensus of these annotations (majority vote for each criterion). We additionally extract bounding box annotations automatically from the 1933 segmentation masks and augment the dataset with these labels Murali et al (2023b).
@@ 3 Dataset
Endoscapes-BBox201 120 41 40 201 36694 12372 9519 58585 Bounding Box 1212 409 312 1933
@@
Endoscapes-CVS201. A collection of 58813 frames from the dissection phase (defined as in Twinanda et al (2016)) of the 201 LC videos. Of these 58813 frames, 11090 frames (1 frame every 5 seconds) are annotated with CVS by three experts, where CVS labels are a collection of three binary image-level annotations indicating achievement of each of the three CVS criteria: namely, C1 - Two Structures, C2 - Hepatocystic Triangle Dissection, and C3 - Cystic Plate, as formalized in Mascagni et al (2020). We split the videos into 120 training, 41 validation, and 40 test, yielding 36694 training frames, 12372 validation frames, and 9747 testing frames; of these, 6970 training frames, 2331 validation frames, and 1799 testing frames contain CVS annotations. To generate the splits, we employ stratified random sampling, using video-level CVS achievement for stratification; splitting the dataset by video (by patient) follows common practice in surgical video analysis Twinanda et al (2016); Zisimopoulos et al (2018); Nwoye and Padoy (2022). These same dataset splits were also used in Alapatt et al (2021); Murali et al (2023b, a); Alapatt et al (2023), thereby facilitating comparisons.
@@
Endoscapes-Seg50. A subset of Endoscapes-Seg201 containing 14940 frames from 50 videos (a subset of the 201 videos), of which 493 (1 frame every 30 seconds) are annotated with segmentation masks. We select the 50 videos by selecting 30, 10, and 10 videos from the aforementioned 120 training, 41 validation, and 40 test videos respectively, employing stratified random sampling based on video-level CVS achievement. This ensures that Endoscapes-Seg50 can be used in conjunction with Endoscapes-BBox201 and Endoscapes-CVS201 to train mixed-supervision models for object detection, segmentation, and CVS prediction. Following the splitting, we have 10380 training, 2310 validation, and 2250 testing frames of which 343, 76, and 74 respectively contain segmentation masks. It should be noted that the 30 training videos of Endoscapes-Seg50 correspond to the third 25% subset fold of Endoscapes-[CVS/Seg]201 that we use for low-label experiments in Alapatt et al (2023); Ramesh et al (2023).
@@
Model Detection mAP@[0.5:0.95] |normal-||| Segmentation mAP@[0.5:0.95]
@@ 4.1 Object Detection
We present the performance of three different object detectors on the Endoscapes2023 test set in Table 2: Faster-RCNN, Cascade-RCNN, and Deformable-DETR, originally reported in Murali et al (2023b). We initialize all detectors using COCO-pretrained weights. Then, for Faster-RCNN and Cascade-RCNN, we train for 20 epochs, while for Deformable-DETR, we train for 30 epochs, following the COCO learning schedule for all detectors. Following most literature in object detection, we report mAP@[0.5:0.95] and breakdown performance per object class.
@@ 4.2 Segmentation
For the two segmentation tasks, we report two performance numbers for each model: performance when trained and tested on Endoscapes-Seg50, as well as the performance of models trained on Endoscapes-Seg201 as in Alapatt et al (2021); Murali et al (2023b, a). This latter performance number represents a ceiling performance that can inform and drive the development of label-efficient segmentation methods. Table 4 shows the performance of Mask-RCNN, Cascade-Mask-RCNN, and Mask2Former for instance segmentation, using segmentation mAP@[0.5:0.95] as the metric. Then, we break down per-class performance in Table 3.
Table 7: CVS Prediction Performance Benchmark on the test set of Endoscapes-CVS201. Models are trained in three different settings, arranged from most label-efficient to least label-efficient. Asterisks (*) indicate spatiotemporal methods. Performances for the first two settings (Endoscapes-CVS201 only and Endoscapes-BBox201 + Endoscapes-CVS201) are cited from Murali et al (2023b, a). Meanwhile, to represent the last experimental setting from Murali et al (2023b, a) that uses Endoscapes-Seg201, we rerun all models using Endoscapes-Seg50 to train the underlying Mask-RCNN model. The performance drop from the results cited in Murali et al (2023b) can be attributed to the use of far fewer segmentation masks during training.
@@ 4.3 CVS Prediction
Finally, we present baseline results for CVS prediction. Murali et al (2023b) contains a detailed exploration of the various possible experimental settings for CVS prediction, along with a benchmark of several single-frame approaches for CVS prediction that can make use of bounding box/segmentation labels if available. Meanwhile, Murali et al (2023a) explores spatiotemporal methods for CVS prediction, again considering the same experimental settings, while Alapatt et al (2023) explores label-efficient CVS prediction. We summarize the three experimental settings below and present results for Endoscapes2023, quoting Murali et al (2023b)Murali et al (2023a), and Alapatt et al (2023) where applicable and rerunning the methods when the dataset used diverges from Endoscapes2023.
The three experimental settings for CVS prediction are: (1) using Endoscapes-CVS201 alone, thereby using only relatively ‘cheap’ CVS annotations, (2) using Endoscapes-CVS201 in addition to Endoscapes-BBox201, and (3) using Endoscapes-CVS201 in addition to Endoscapes-Seg50. For the last setting, as for the segmentation benchmark, we rerun the methods from Murali et al (2023b) and Murali et al (2023a) using only Endoscapes-Seg50 as opposed to Endoscapes-Seg201. We additionally present the original results using Endoscapes-Seg201 + Endoscapes-CVS201 from Murali et al (2023b, a) to illustrate ceiling performance.
@@ 4.4.1 Endoscapes-CVS201 Only
ResNet50. We show the results of a vanilla ResNet50 classifier with the final fully-connected layer modified to output three scores corresponding to the CVS criteria (from Murali et al (2023b)).
ResNet50 with Reconstruction. The same as the above with an added auxiliary image reconstruction objective (from Murali et al (2023b)).
ResNet50-MoCov2. Same ResNet50 classifier but initialized using a MoCov2-pretrained ResNet50 backbone (corresponds to the best performing initialization from Alapatt et al (2023).
@@ 4.4.2 Endoscapes-BBox201 + Endoscapes-CVS201
LayoutCVS. A Faster-RCNN object detector is trained using the subset of images with bounding box annotations. Then, for CVS prediction, the trained, frozen object detector is run on an image, the output boxes are reorganized into a $C$-dimensional, pixel-wise layout, and this layout is passed to a ResNet18 classifier to predict the three CVS criteria. $C$ refers to the number of object classes in the dataset, which in this case is $6$. Performance reported from Murali et al (2023b).
DeepCVS. The same as LayoutCVS, but the layout is concatenated with the original image prior to classification. Performance reported from Murali et al (2023b).
ResNet50-DetInit. A ResNet50 classifier with the backbone weights initialized from the trained object detector. Performance reported from Murali et al (2023b).
LG-CVS. The method proposed in Murali et al (2023b), which extracts per-object features based on the detected objects and builds a latent graph with object semantic and visual properties as well as inter-object relations. This latent graph is then used to predict CVS. Performance reported from Murali et al (2023b).
DeepCVS-Temporal*. A spatiotemporal extension of DeepCVS proposed in Murali et al (2023a) that computes ResNet18 embeddings of a sequence of $T$ frames and processes with a Transformer to predict CVS. Performance when setting $T=10$ is reported from Murali et al (2023a).
STRG*. A spatiotemporal graph-based approach originally proposed in Wang and Gupta (2018) for fine-grained action recognition and adapted for CVS prediction in Murali et al (2023a). Performance when using a clip size of $10$ is reported from Murali et al (2023a).
SV2LSTG*. The method proposed in Murali et al (2023a) entitled “Encoding Surgical Videos as Latent SpatioTemporal Graphs”, where the LG-CVS Murali et al (2023b) encoder is used to extract a series of single frame graphs that are then linked temporally and used to predict CVS with a graph neural network. Performance when using a clip size of $10$ is reported from Murali et al (2023a).
@@ 4.4.3 Endoscapes-Seg50 + Endoscapes-CVS201
We also report the results when running all the methods from Section 4.4.2 using Endoscapes-Seg50 instead of Endoscapes-BBox201; for all methods, we train a Mask-RCNN instance segmentation model in place of the Faster-RCNN object detector. This represents the most label-intensive experimental setting. We also report the original results of Murali et al (2023b, a), where each model is trained on Endoscapes-Seg201 + Endoscapes-CVS201, to represent ceiling performance.
@@ 4.4.4 Low-Label Settings
Lastly, in Table 8 we show selected results on two lower training data budgets of Endoscapes-CVS201: 12.5% and 25%, reporting results from Alapatt et al (2023). These experiments, and the 12.5% data budget in particular, represent the most label-efficient evaluation setting, using a small subset of annotations and requiring neither bounding box nor segmentation annotations, both of which are more expensive to collect than CVS labels. For each budget, we sample three different sets of videos (15/30 videos respectively) from the train set of Endoscapes-CVS201. The results using the entire train set of Endoscapes-CVS201 are shown in the first section of Table 7 (Endoscapes-CVS201 Only), and represent a ceiling performance for these experiments. We release these splits along with the remainder of the dataset to facilitate comparisons with the methods presented here.

### v3

@@ 1 Introduction
Automatic assessment of the critical view of safety (CVS) is an important problem in surgical data science that has been gaining interest in recent years [13, 5, 7, 9, 3, 12, 8]. Unlike many surgical data science tasks that have been studied thus far (e.g. phase recognition, tool detection/segmentation), CVS assessment is particularly challenging because it relies on accurate and precise identification of fine-grained anatomical structures and concepts. To enable research into this task, we present Endoscapes, a dataset of laparoscopic cholecystectomy videos with annotations specifically targeted at tackling automated CVS assessment: namely, frame-level CVS annotations as well as segmentation/bounding boxes of critical anatomical structures/regions relevant for CVS assessment (such as the hepatocystic triangle, cystic artery/duct, and cystic plate).
In this report, we begin by describing the development of this dataset, which represents the culmination of several works [7, 1, 9]. Then, we provide a detailed enumeration of the contents of Endoscapes2023, released alongside this report, which comprises three sub-datasets categorized by annotation type and number of videos: Endoscapes-CVS201, Endoscapes-BBox201, and Endoscapes-Seg50. Finally, we present benchmarks for three different tasks (object detection, instance segmentation, and CVS prediction).
@@ 2 Context
Automated CVS assessment was first tackled by our prior work DeepCVS [7], where we present a method for joint anatomical segmentation and CVS prediction. For this initial work, we hand-picked 2854 images where we deemed CVS assessable (in the anterior view) from 201 laparoscopic cholecystectomy (LC) videos, then annotated these images with each of the three CVS criteria (three binary image-level annotations), based on the protocol established in [5]. We also selected a subset of 402 images ($\sim$20%), balancing optimal CVS achievement (3/3 criteria achieved) and suboptimal CVS achievement (0-2/3 criteria achieved) and further annotated them with segmentation masks. This initial dataset, while instrumental in enabling a proof-of-concept study, was characterized by two critical limitations: (1) the frame selection process introduces bias during both training and evaluation, and (2) the dataset is very small, with a test set of only 571 images for CVS assessment and 80 images for segmentation.
Recognizing these issues, in a follow-up work [1], we collected a vastly expanded dataset from the same 201 videos. Here, rather than hand-selecting frames from the videos, we simply extract frames at 1fps from the region of video where CVS is adjudged to be evaluable [^4] This region, which is a subset of the dissection phase, begins when “any of the 3 CVS criteria is considered evaluable” and ends “when the first clip is applied on the cystic duct or the cystic artery”, as defined in [6]., thereby eliminating selection bias. For these frames, we then annotate segmentation (5 anatomical structures and 1 tool class) and CVS at regular intervals: 1 frame every 30 seconds for segmentation [1] and 1 frame every 5 seconds for CVS [9]. Removing the manual frame selection process introduced numerous issues during annotation, to which end we developed a more robust annotation protocol for both segmentation and CVS assessment [6]. Finally, following this updated protocol and incorporating several rounds of review to ensure optimal consistency, we annotated 1933 total images with segmentation masks and 11090 images with CVS annotations from the 201 LC videos. The CVS annotations in particular were done independently by three expert clinicians to reduce bias; for evaluation, we define the ground-truth as the consensus of these annotations (majority vote for each criterion). We additionally extract bounding box annotations automatically from the 1933 segmentation masks and augment the dataset with these labels [9].
@@ 3 Dataset
Endoscapes-BBox201 120 41 40 201 36694 12372 9747 58813 Bounding Box 1212 409 312 1933
@@
Endoscapes-CVS201. A collection of 58813 frames from the dissection phase (defined as in [14]) of the 201 LC videos. Of these 58813 frames, 11090 frames (1 frame every 5 seconds) are annotated with CVS by three experts, where CVS labels are a collection of three binary image-level annotations indicating achievement of each of the three CVS criteria: namely, C1 - Two Structures, C2 - Hepatocystic Triangle Dissection, and C3 - Cystic Plate, as formalized in [5]. Cohen’s kappa for inter-rater agreements for C1, C2, and C3, and the overall CVS assessment (1 if all 3 criteria are achieved, 0 otherwise) were $0.33$, $0.53$, $0.44$, and $0.38$. These Cohen’s kappa values were first computed between each pair of annotators, then averaged across the three possible pairs for each of the three criteria and overall CVS.
We split the videos into 120 training, 41 validation, and 40 test, yielding 36694 training frames, 12372 validation frames, and 9747 testing frames; of these, 6960 training frames, 2331 validation frames, and 1799 testing frames contain CVS annotations. To generate the splits, we employ stratified random sampling, using video-level CVS achievement for stratification; splitting the dataset by video (by patient) follows common practice in surgical video analysis [14, 16, 10]. These same dataset splits were also used in [1, 9, 8, 2], thereby facilitating comparisons.
@@
Endoscapes-Seg50. A subset of Endoscapes-Seg201 containing 14940 frames from 50 videos (a subset of the 201 videos), of which 493 (1 frame every 30 seconds) are annotated with segmentation masks. We select the 50 videos by selecting 30, 10, and 10 videos from the aforementioned 120 training, 41 validation, and 40 test videos respectively, employing stratified random sampling based on video-level CVS achievement. This ensures that Endoscapes-Seg50 can be used in conjunction with Endoscapes-BBox201 and Endoscapes-CVS201 to train mixed-supervision models for object detection, segmentation, and CVS prediction. Following the splitting, we have 10380 training, 2310 validation, and 2250 testing frames of which 343, 76, and 74 respectively contain segmentation masks. It should be noted that the 30 training videos of Endoscapes-Seg50 correspond to the third 25% subset fold of Endoscapes-[CVS/Seg]201 that we use for low-label experiments in [2, 11].
@@
Model Detection mAP@[0.5:0.95] |||| Segmentation mAP@[0.5:0.95]
@@ 4.1 Object Detection
We present the performance of three different object detectors on the Endoscapes2023 test set in Table 2: Faster-RCNN, Cascade-RCNN, and Deformable-DETR, originally reported in [9]. We initialize all detectors using COCO-pretrained weights. Then, for Faster-RCNN and Cascade-RCNN, we train for 20 epochs, while for Deformable-DETR, we train for 30 epochs, following the COCO learning schedule for all detectors. Following most literature in object detection, we report mAP@[0.5:0.95] and breakdown performance per object class.
@@ 4.2 Segmentation
For the two segmentation tasks, we report two performance numbers for each model: performance when trained and tested on Endoscapes-Seg50, as well as the performance of models trained on Endoscapes-Seg201 as in [1, 9, 8]. This latter performance number represents a ceiling performance that can inform and drive the development of label-efficient segmentation methods. Table 4 shows the performance of Mask-RCNN, Cascade-Mask-RCNN, and Mask2Former for instance segmentation, using segmentation mAP@[0.5:0.95] as the metric. Then, we break down per-class performance in Table 3.
Table 7: CVS Prediction Performance Benchmark on the test set of Endoscapes-CVS201. Models are trained in three different settings, arranged from most label-efficient to least label-efficient. Asterisks (*) indicate spatiotemporal methods. Performances for the first two settings (Endoscapes-CVS201 only and Endoscapes-BBox201 + Endoscapes-CVS201) are cited from [9, 8]. Meanwhile, to represent the last experimental setting from [9, 8] that uses Endoscapes-Seg201, we rerun all models using Endoscapes-Seg50 to train the underlying Mask-RCNN model. The performance drop from the results cited in [9] can be attributed to the use of far fewer segmentation masks during training.
@@ 4.3 CVS Prediction
Finally, we present baseline results for CVS prediction. [9] contains a detailed exploration of the various possible experimental settings for CVS prediction, along with a benchmark of several single-frame approaches for CVS prediction that can make use of bounding box/segmentation labels if available. Meanwhile, [8] explores spatiotemporal methods for CVS prediction, again considering the same experimental settings, while [2] explores label-efficient CVS prediction. We summarize the three experimental settings below and present results for Endoscapes2023, quoting [9][8], and [2] where applicable and rerunning the methods when the dataset used diverges from Endoscapes2023.
The three experimental settings for CVS prediction are: (1) using Endoscapes-CVS201 alone, thereby using only relatively ‘cheap’ CVS annotations, (2) using Endoscapes-CVS201 in addition to Endoscapes-BBox201, and (3) using Endoscapes-CVS201 in addition to Endoscapes-Seg50. For the last setting, as for the segmentation benchmark, we rerun the methods from [9] and [8] using only Endoscapes-Seg50 as opposed to Endoscapes-Seg201. We additionally present the original results using Endoscapes-Seg201 + Endoscapes-CVS201 from [9, 8] to illustrate ceiling performance.
@@ 4.4.1 Endoscapes-CVS201 Only
ResNet50. We show the results of a vanilla ResNet50 classifier with the final fully-connected layer modified to output three scores corresponding to the CVS criteria (from [9]).
ResNet50 with Reconstruction. The same as the above with an added auxiliary image reconstruction objective (from [9]).
ResNet50-MoCov2. Same ResNet50 classifier but initialized using a MoCov2-pretrained ResNet50 backbone (corresponds to the best performing initialization from [2].
@@ 4.4.2 Endoscapes-BBox201 + Endoscapes-CVS201
LayoutCVS. A Faster-RCNN object detector is trained using the subset of images with bounding box annotations. Then, for CVS prediction, the trained, frozen object detector is run on an image, the output boxes are reorganized into a $C$-dimensional, pixel-wise layout, and this layout is passed to a ResNet18 classifier to predict the three CVS criteria. $C$ refers to the number of object classes in the dataset, which in this case is $6$. Performance reported from [9].
DeepCVS. The same as LayoutCVS, but the layout is concatenated with the original image prior to classification. Performance reported from [9].
ResNet50-DetInit. A ResNet50 classifier with the backbone weights initialized from the trained object detector. Performance reported from [9].
LG-CVS. The method proposed in [9], which extracts per-object features based on the detected objects and builds a latent graph with object semantic and visual properties as well as inter-object relations. This latent graph is then used to predict CVS. Performance reported from [9].
DeepCVS-Temporal*. A spatiotemporal extension of DeepCVS proposed in [8] that computes ResNet18 embeddings of a sequence of $T$ frames and processes with a Transformer to predict CVS. Performance when setting $T=10$ is reported from [8].
STRG*. A spatiotemporal graph-based approach originally proposed in [15] for fine-grained action recognition and adapted for CVS prediction in [8]. Performance when using a clip size of $10$ is reported from [8].
SV2LSTG*. The method proposed in [8] entitled “Encoding Surgical Videos as Latent SpatioTemporal Graphs”, where the LG-CVS [9] encoder is used to extract a series of single frame graphs that are then linked temporally and used to predict CVS with a graph neural network. Performance when using a clip size of $10$ is reported from [8].
@@ 4.4.3 Endoscapes-Seg50 + Endoscapes-CVS201
We also report the results when running all the methods from Section 4.4.2 using Endoscapes-Seg50 instead of Endoscapes-BBox201; for all methods, we train a Mask-RCNN instance segmentation model in place of the Faster-RCNN object detector. This represents the most label-intensive experimental setting. We also report the original results of [9, 8], where each model is trained on Endoscapes-Seg201 + Endoscapes-CVS201, to represent ceiling performance.
@@ 4.4.4 Low-Label Settings
Lastly, in Table 8 we show selected results on two lower training data budgets of Endoscapes-CVS201: 12.5% and 25%, reporting results from [2]. These experiments, and the 12.5% data budget in particular, represent the most label-efficient evaluation setting, using a small subset of annotations and requiring neither bounding box nor segmentation annotations, both of which are more expensive to collect than CVS labels. For each budget, we sample three different sets of videos (15/30 videos respectively) from the train set of Endoscapes-CVS201. The results using the entire train set of Endoscapes-CVS201 are shown in the first section of Table 7 (Endoscapes-CVS201 Only), and represent a ceiling performance for these experiments. We release these splits along with the remainder of the dataset to facilitate comparisons with the methods presented here.
@@
4.5 Training Details
We follow the implementation and training settings of [9, 2] using the mmdetection framework [4] for all reported methods. CVS prediction models are trained with a weighted binary cross entropy loss, where the weight for each criterion is computed by inverse frequency balancing. All object detection and instance segmentation models are initialized with COCO-pretrained weights and finetuned.